# Interpretable Retinal Disease Prediction Using Biology-Informed Heterogeneous Graph Representations

### v1

@@ V Interpretability Analysis
Figure 6: Comparison of different explainability tools for the DR staging task on OCTA images. GradCAM and integrated gradients are state-of-the-art approaches that previous works have applied to DR staging. Sample a) describes a sample with clinical diagnosis NPDR that the CNN and graph classification model correctly identified. Sample b) describes a sample with clinical diagnosis PDR that the CNN and graph classification model correctly identified. Compared to the traditional spatial explanations, our framework (rightmost column) allows the precise localization of critical vessel segments and intercapillary areas.
To illustrate the human-understandable explanations of our interpretability framework, we analyze two samples from the test set. Figure 5 a) and b) depict the spatial and feature-wise explanations for two samples that were correctly classified as NPDR and PDR.
V-A Qualitative Results
In both samples, we highlight the twenty most important nodes in their associated image region. In a), we show that large intercapillary areas and some small capillaries are most influential to the model’s decision. Conversely, in Figure 5 b), a combination of large and small vessels, intercapillary areas, and the FAZ are identified as the most important nodes.
Beyond spatial explanations, our method enables us to attribute the elevated importance of specific nodes to high-level concepts, such as vessel curvature or intercapillary area size. The bar plots in Figure 5 indicate which features of the highlighted nodes are decisive for the predictions. The arrows in the right-most column of Figure 5 indicate how the respective feature deviates from the average.
In sample a), the increased number of neighboring intercapillary areas and the larger perimeter of the intercapillary areas are the most influential characteristics. This observation aligns with earlier studies reporting a relation between the size of the largest intercapillary areas and DR progression [39]. The study points out that the emergence of large intercapillary areas is caused by vascular dropout, which results in the merging of two areas. The number of neighboring intercapillary areas is, therefore, also expected to increase. Furthermore, the reduced number of neighbors at the smaller branching point and the increased deviation of the roundness are important vessel characteristics.
In sample b), the high curvature and the diameter variability are the predominant vessel characteristics for the prediction. These characteristics can be related to intraretinal microvascular abnormalities (IRMA) such as vessel dilations and to the neovascularization of tiny vessel loops [55]. The most influential properties of the intercapillary areas are an increased area size and a large number of neighboring areas. Interestingly, the FAZ also has a strong influence on the prediction through a decrease in solidity, defined as its area divided by the area of its convex hull. This explanation is also related to previous studies reporting that peripheral disruptions of the FAZ, which can result in a decrease of solidity, are linked to DR [56]. Numerous studies identified the FAZ circularity, a closely related concept, as predictive for DR [41].
V-B Comparison to CNN Explanations
We adopt explainability approaches previously employed by Ryu et al. [10] and Heisler et al. [9] and use grad-CAM to create model explanations. Additionally, we utilize integrated gradients, initially proposed in the DR classification setting [43], to highlight finer structures than grad-CAM. Moreover, we experiment with an alternative integrated gradients approach for CNNs, using a baseline input that accounts for the expected lower image intensity around the FAZ. This is achieved by averaging the normalized image intensities across the training set, resulting in the elimination of highlighted areas within the FAZ caused by noise.
@@ V-B Comparison to CNN Explanations
Grad-CAM explanations encompass a broad area, spanning multiple vessels and intercapillary areas. Therefore, it is infeasible to identify the exact entities that influence the prediction. Furthermore, it is impossible to identify which characteristics of the respective image regions cause their elevated importance. An analysis of the average grad-CAM explanation revealed a consistent focus on the FAZ, which is also visible in Figure 6.
Figure 7: Quantitative evaluation of a), the most important characteristics for correct identification of the PDR stage in the test set, using our proposed graph classification model. The red and blue colors represent characteristics of the vessels and intercapillary areas, respectively. b) Violin plots of the average curvature across all vessel segments in a graph according to their clinical label. c) Violin plots of the average area in no. of pixels across all intercapillary areas in a graph according to their clinical label.
In contrast, explanations using integrated gradients are more precise in location, sometimes even highlighting individual vessel segments. However, in most cases, only a few connected pixels are highlighted, and many of these clusters are dispersed across the image. We observe, across all images, that a significant portion of the highlighted pixels is in the FAZ without any notable deviations between samples. Adopting an average OCTA image as the baseline sample mitigates this behavior. However, this approach leads to vessels close to the FAZ being highlighted, which are not highlighted in the zero image approach. This effect arises from the newly introduced stronger contrast to the average pixel intensity in the FAZ region. Baseline images strongly affect the created explanations, and ideally, multiple baselines are considered to obtain in-depth explanations [45].
V-C Global Feature Importance
Beyond examining model reasoning on single samples (Figure 5), we investigate global trends of influential features for decision-making. In Figure 7 a), we show the relative importance of the most influential characteristics for correctly classifying PDR. We identify vessel curvature and the size of the intercapillary areas as the most important features for the respective node types. Interestingly, the FAZ’s properties are less important than the vessel and intercapillary area characteristics. Moreover, the distributions of the average curvature (7 b)) and average intercapillary area size (7 c)) show that globally, an increase of curvature and area size is directly associated with the PDR and NPDR stage. Previous studies found that vessel tortuosity is linked to the NPDR, but interestingly not the PDR stage [57]. This distinction may arise from the difference in the average curvature, where vessel segments are weighted equally, and the vessel tortuosity, where vessels are weighted according to their length. The increase in intercapillary area size was linked to DR progression earlier [39]. Moreover, in the case of the loss of separating vessels, it is directly related to VD, which has been associated with DR in numerous studies [41].
VI Discussion
@@
In this work, we propose an interpretable method for clinical DR grading in OCTA images. We establish a specialized heterogeneous graph representation that models the particularities of the retinal vasculature. Building on this novel data representation, we develop an efficient GNN model that solves DR staging as a graph classification task, surpassing image-based methods in performance. Finally, we introduce an interpretability tool providing fine-grained explanations for our predictions using spatial explanations that highlight critical structures and attribute their importance to specific characteristics.
@@
Our classification approach provides a notable step towards interpretable disease classification of OCTA images, which is crucial for adapting discriminative models in clinical practice. While our method improves existing explainability frameworks, we have not created a perfectly interpretable classification pipeline. Our improvements concern perspective and mathematical interpretability. Perspective interpretability is improved by the fine-grained spatial explanations that allow us to precisely visualize important vessels and intercapillary areas, which has been impossible for CNNs and vision transformers. Moreover, our attributions consist of human-interpretable concepts that are commonly associated with disease progression. By introducing a data representation based on the retinal vasculature, we restrict our model to these anatomical constraints in decision-making. We improve mathematical interpretability by limiting learnable parameters and network layers [58]. Nevertheless, our model consists of multiple layers with activations and numerous features, which allows for non-trivial feature interaction. Therefore, the study of feature interaction, which has been previously done for MLPs [59], is a necessary future work.
@@
Although our approach avoids using the largely uninterpretable CNN and transformer models for the classification stage, it still relies on these architectures to generate high-quality segmentations. Utilizing this class of networks in the data representation stage provides some advantages compared to the usage at the later stage. The first advantage is the enhanced interpretability of the algorithmic outcomes using the abstracted representation. The second advantage is that the vision models can be trained on existing large-scale synthetic datasets [14] compared to the disease classification task, where high-quality data is still limited. A limitation persists in the susceptibility of the graph representation to segmentation errors. However, we argue that single-stage CNN or vision transformer models would implicitly create internal representations similar to a segmentation and, therefore, may be similarly affected. Moreover, segmentation errors are easier to detect than reasoning about erroneous internal representations in CNNs and vision transformers. Beyond image classification, we believe that our representation is a foundation for future applications, such as link prediction for correcting erroneous segmentation and anomaly detection of intercapillary areas or vessel segments, considering their properties and neighborhoods. Finally, disease prediction through graph classification, as shown in this work, can be extended to a diverse set of clinically relevant diseases.

### v2

@@ 5.1 Feature Relation to Microvascular Pathologies
Figure 7: Comparison of different explainability tools for the DR staging task on OCTA images. GradCAM and integrated gradients are state-of-the-art approaches that previous works have applied to DR staging. Sample a) describes a sample with clinical diagnosis NPDR that the CNN and graph classification model both classify as PDR. Sample b) describes a sample with clinical diagnosis PDR that the CNN and graph classification model both also identify as PDR. Compared to traditional spatial explanations for vision models, our framework (bottom row) enables the precise localization of critical vessel segments with shape and intensity that indicate intraretinal microvascular abnormalities and perfused microaneurysms. Intercapillary areas indicate image areas where our model sees reduced capillary density.
5.2 Examples for spatial explanations
To illustrate that (1) our method links pathological microvascular changes to the stage prediction task, and (2) our method allows traceable prediction results, we show explanations for predicted DR stages. Our method allows the spatially exact identification of the most important vessel segments and intercapillary areas for the DR stage prediction task. Figure 7 shows two examples where the most important vessel segments and intercapillary areas are highlighted for two patients suffering from DR. The precise spatial attribution is in stark contrast to common CNN explanations, which often provide broad and scattered explanations that do not allow for a precise attribution to abnormal vessels, areas of capillary dropout, or microaneurysms, guiding the prediction. We adopt explainability approaches previously employed by Ryu et al. [ryu2021deep] and Heisler et al. [heisler2020ensemble] and use grad-CAM to create model explanations. The grad-CAM explanations are focused on the center of the FAZ, which is unrelated to DR disease progression. Similarly, a majority of the area highlighted by integrated gradients, combined with the vision models, highlights areas within the FAZ. For the subject in 7 a) where the model prediction on OCTA deviates from the clinical fundus grading, both vision explanations do not provide an interpretable explanation for the prediction of the further progressed DR stage. In contrast, our methods clearly highlight vessel segments that resemble abnormal structures that guide the model to predict late-stage DR, deviating from the fundus-based diagnosis.
Figure 8: Explanations for model predictions and explanations of our framework with predicted stages agreeing to color fundus staging (b), c)), and an example with diverging prediction (a)). Red overlay indicates contribution of interpcaillary areas, FAZ (center column), or vessel segments (right column) to the predicted stage. Boxes below the images indicate which features are most important to the prediction, and how they diverge from the norm.
Beyond the spatial explanations, our method enables us to attribute the elevated importance of specific nodes to high-level concepts, such as vessel tortuosity or intercapillary area size, that are directly linked to disease progression. We display three different examples: Figure 8 a) shows an example where the clinical label is no DR while the OCTA-based prediction as well as the OCTA expert label indicate PDR, Figure 8 b) displays an image that was classified as PDR with clinical label PDR, and Figure 8 c) displays an image that was classified as NPDR with clinical label NPDR. We indicate the most important characteristics and their deviation from the norm for exemplary nodes. Additionally, we provide full interactive explanations in our code repository, which allows us to see the influence of intercapillary areas, the FAZ, vessel segments, and their characteristics on the predicted DR stage.
Example a) in Figure 8 shows the OCTA image of a subject with no DR following standard fundus examination. The OCTA image was graded with the PDR stage by an expert ophthalmologist. Our model aligns with the OCTA expert analysis, highlighting the intercapillary areas around the FAZ as highly important, due to their increased size (e.g., minor axis length and major axis length) and their low average image intensity. The eccentricity of the FAZ also contributed to the stage prediction. Moreover, specific vessel segments contribute to the PDR diagnosis, such as the bright intensity terminal vessel emphasized by the red arrow. In the prediction, the size (area and average radius), shape abnormality (increased radius variability), hyperintensity, and the terminal position are the most important characteristics.
Example b) in Figure 8 shows an example where the clinical stage is PDR, aligning with the severity of the microvascular changes in the OCTA image. While the FAZ does not indicate the PDR stage, the large size and low intensity in the capillary areas are indicative of the PDR stage. Moreover, the model identifies multiple abnormally shaped vessels, such as the vessel segment indicated by the red arrow, which exhibits increased size (area + radius), atypical shape characteristics (increased radius variability), and unusually high variation of intensity within the segment.
Finally, example c) in Figure 8 shows the OCTA image of a subject with clinical label NPDR and model prediction NPDR. The decreased solidity of the FAZ and the slightly increased (+0.3 SD) major axis length contribute to the NPDR prediction. Moreover, some intercapillary areas contribute to the DR prediction through increased size and decreased intensity. The vessel segment indicated by the red arrow, which shows strong similarities to a microaneurysm, strongly contributes to the NPDR diagnosis through size (area and radius), shape (radius variability), hyperintensity, and terminal position.
@@ 5.2 Examples for spatial explanations
5.3 Comparison to Other Graph-Based Explanations
Random forest and support vector machine classifiers on aggregated node embeddings come with the disadvantage of losing spatial information for explainability. The aggregation of node information followed by an ensemble of decision trees in random forests does not allow spatial hinting towards critical regions.
Homogeneous graph representations do not allow for a direct attribution towards clinically well-studied microvascular changes. Using vessel graphs alone does not allow direct attribution to regions of ischemia and capillary dropout, which are hallmarks of DR. Using intercapillary areas alone does not allow attribution to intraretinal microvascular abnormalities, which are the second pillar for the analysis of DR in OCTA images. Only the combination of both representations provides the capability to attribute the classification directly to the microvascular changes that comprise dropout, abnormalities, and new-growth of vessels.
6 Discussion
@@
In this work, we propose an interpretable method for DR staging support using OCTA images. We establish a specialized heterogeneous graph representation that models the particularities of the retinal vasculature. Building on this novel data representation, we develop an efficient GNN model that solves DR staging as a graph classification task. Finally, we introduce an interpretability tool providing fine-grained explanations for our predictions using spatial explanations that highlight critical structures and attribute their importance to specific characteristics.
@@
Our staging approach provides a notable step towards interpretable disease staging of OCTA images, which is crucial for adapting discriminative models in clinical practice. While our method improves existing explainability frameworks, we have not created a perfectly interpretable classification pipeline. Our improvements concern perspective and mathematical interpretability. Perspective interpretability is improved by the fine-grained spatial explanations that allow us to precisely visualize important vessels and intercapillary areas, which has been impossible for CNNs and vision transformers. Moreover, our attributions consist of human-interpretable concepts that are commonly associated with disease progression. By introducing a data representation based on the retinal vasculature, we restrict our model to these anatomical constraints in decision-making. We improve mathematical interpretability by limiting learnable parameters and network layers [barcelo2020model]. Nevertheless, our model consists of multiple layers with activations and numerous features, which allows for non-trivial feature interaction. Therefore, the study of feature interaction, which has been previously done for MLPs [tsang2017detecting], is a necessary future work.
@@
Although our approach avoids using the largely uninterpretable CNN and transformer models for the classification stage, it still relies on these architectures to generate high-quality segmentations. Utilizing this class of networks in the data representation stage provides some advantages compared to the usage at a later stage. The first advantage is the enhanced interpretability of the algorithmic outcomes using the abstracted representation. The second advantage is that the vision models can be trained on existing large-scale synthetic datasets [kreitner2023detailed] compared to the disease classification task, where high-quality data is still limited. A limitation persists in the susceptibility of the graph representation to segmentation errors. However, we argue that single-stage CNN or vision transformer models would implicitly create internal representations similar to a segmentation and, therefore, may be similarly affected. Moreover, segmentation errors are easier to detect than reasoning about erroneous internal representations in CNNs and vision transformers.
Beyond image classification, we believe that our representation is a foundation for future applications, such as link prediction for correcting erroneous segmentation and node classification of intercapillary areas or vessel segments, e.g., for the identification of microaneurysms, intraretinal microvascular abnormalities, and neovascularization.
Appendix A Acknowledgements
A.1 Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
A.2 Generative AI
We used large language models at the sentence level to improve formulations and correct spelling and grammar.